OpenStudy (anonymous):

Express the product of x+2/2 and 4x+20/X^2+6x+8

OpenStudy (solomonzelman):

\(\LARGE\color{blue }{ \rm \frac{x+2}{x} \times \frac{4x+20}{x^2+6x+8} }\)

OpenStudy (solomonzelman):

\(\LARGE\color{blue }{ \rm \frac{(x+2)4(x+5)}{x(x+2)(x+4)} }\)

OpenStudy (solomonzelman):

\(\LARGE\color{blue }{ \rm \frac{4(x+5)}{x(x+4)} }\)

OpenStudy (anonymous):

That's the answer?

OpenStudy (solomonzelman):

yes, you can expand the top and the bottom, but the answer is the same.

OpenStudy (anonymous):

so its simplified already, right?

OpenStudy (solomonzelman):

it is simplified. What you can you is expand the top and bottom, but that doesn't make a difference.

OpenStudy (anonymous):

Ok Thank You.

OpenStudy (solomonzelman):

Anytime !